OpenStudy (anonymous):

Find slope of line passing through the pair of points (3,6) (-6,4)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (4-6)/(-6-3) m = (-2)/(-9) m = 2/9 So the slope of the line that passes through the two points is 2/9

OpenStudy (amistre64):

subtract one point from the other and stack y/x

OpenStudy (amistre64):

(3,6) -(-6,4) ------ 9,2 ; slope = 2/9

OpenStudy (anonymous):

I thought slope was y=mx+b.

OpenStudy (amistre64):

slope is NOT the equation of a line; it is PART OF the equation

OpenStudy (amistre64):

y = slope x + b

OpenStudy (anonymous):

OH okayy thanks